OpenStudy (anonymous):

a set of real numberssatifying the fiven inequality is one or more intervals abvalue x+pie/2 absvalue>1

OpenStudy (across):

Find a range for x such that\[|x+\frac{\pi}{2}|>1\]Is that what you're asking?

OpenStudy (anonymous):

I have to find the real nuymber satifying the give inequality

OpenStudy (across):

The inequality I posted above, you mean?

OpenStudy (anonymous):

yes

OpenStudy (across):

All you have to do is solve for x in the equation: \[x+\frac{1}{2}\pi=1\]Simply enough: \[x=1-\frac{1}{2}\pi\]\[x=\frac{2-\pi}{2}\]But since we know that we're taking the absolute value, we find that \[(-\infty, \frac{2-\pi}{2})∪(\frac{2-\pi}{2}, \infty)\]is our range.

OpenStudy (anonymous):

I got the same thing thank you